A solenoid has 80 cm length, number of loop is 4 and magnetic field inside it

        is 1.2 x 10 – 5T.  Find the current passing through the each loop of wire.



1
Expert's answer
2022-01-26T07:41:45-0500

"B=\u03bc_onI"

"μ_o=4\pi×10^-7"

"n=\frac{4}{0.8}=5"

"B=1.2\u00d710^{-5}T"

"I=\frac{B}{μ_on}"

"I=\frac{1.2×10^{-5}}{4\pi×10^-7×5}=1.9A"